OpenStudy (ibrafacts):

If 5k + 8=-7 + 2k, what is the value of 3k?

OpenStudy (ibrafacts):

how would I approach this problem?

OpenStudy (janu16):

First: solve the "k" variable. combine "like" terms - subtract 2k from both sides

OpenStudy (ibrafacts):

oh

OpenStudy (janu16):

thats what you would have 5K - 2k+8 = -7+2k - 2k 5K - 2k+8 = -7 3k+8 = - 7

OpenStudy (janu16):

than subtract 8 from both sides...

OpenStudy (ibrafacts):

k=-5

OpenStudy (janu16):

3k+8 -8 = - 7 - 8 3k = - 7 - 8 3k = - 15 than divide both sides by 3k

OpenStudy (janu16):

yes but wait

OpenStudy (janu16):

3k/3 = - 15/3 k = -15/3 k = - 5 so we have -5 right?

OpenStudy (ibrafacts):

do i plug -5 into 3k?

OpenStudy (janu16):

Now, find the value of 3k. Replace (-5) with the k" variable in 3k.

OpenStudy (janu16):

what do you get?

OpenStudy (ibrafacts):

-15?

OpenStudy (janu16):

yes good

OpenStudy (ibrafacts):

thank you

OpenStudy (janu16):

yw